OpenStudy (zenmo):

Find the exact area of the surface obtained by rotating the curve about the x-axis.

OpenStudy (zenmo):

\[x = \frac{ 1 }{ 3 }(y^2+2)^{3/2}, 4 \le y \le 5\]

myininaya (myininaya):

you use \[\int\limits 2 \pi y ds \\ \text{ where } ds=\sqrt{1+(y')^2} dx , a \le x \le b \\ \text{ or } \\ \text{ use } ds=\sqrt{1+(x')^2} dy , c \le y \le d \\ \\ \text{ by the way when I say } y' \text{ I mean } \frac{dy}{dx} \\ \text{ and when I said } x' \text{ I mean } \frac{dx}{dy}\]

myininaya (myininaya):

I think I prefer the second ds

OpenStudy (zenmo):

\[x'=\frac{ 1 }{ 3 }[\frac{ 3 }{ 2 }[y^2+2]^{1/2}*(2y)]\] Did I find the x' correctly?

myininaya (myininaya):

looks good can be simplified a bit

OpenStudy (zenmo):

\[x ' = y \sqrt{y^2+2}\]

myininaya (myininaya):

\[x'=y(y^2+1)^\frac{1}{2}\] yep yep

myininaya (myininaya):

and (x')^2=?

OpenStudy (zenmo):

\[y^2(y^2+2)\]

myininaya (myininaya):

you probably got through all of this but got scared of the square root part I will give you a hint 1+(x')^2 ends up being a perfect square

myininaya (myininaya):

good you used your x' I made a type-o in my above :p \[x'=y(y^2+2)^\frac{1}{2} \\ (x')^2=y^2(y^2+2) \\ (x')^2=y^4+2y^2 \\ (x')^2+1=y^4+2y^2+1\]

myininaya (myininaya):

\[(x')^2+1=(y^2+1)^2 \]

myininaya (myininaya):

\[\sqrt{(x')^2+1}=\sqrt{(y^2+1)^2}=y^2+1 \text{ since } y^2+1 \text{ is always positve }\]

OpenStudy (zenmo):

\[S = 2\pi \int\limits_{4}^{5}y \sqrt{1+y^4+2y^2}dy\]

OpenStudy (zenmo):

\[S = 2\pi \int\limits_{4}^{5}y(y^2+1)dy=2\pi \int\limits_{4}^{5}(y^3+y)dy\] Like that?

myininaya (myininaya):

yes

OpenStudy (zenmo):

Okay, i got the problem now, just needed help setting the square root up and using x' instead of y'. Thanks :D

myininaya (myininaya):

np